\begin{thebibliography}{}
\expandafter\ifx\csname natexlab\endcsname\relax\def\natexlab#1{#1}\fi

\bibitem[{Brody and Elhadad(2010)}]{Brody2010}
Samuel Brody and Noemie Elhadad. 2010.
\newblock An unsupervised aspect-sentiment model for online reviews.
\newblock In {\em NAACL '10\/}. pages 804--812.

\bibitem[{Chen and Liu(2014)}]{ChenLiu2014ICML}
Zhiyuan Chen and Bing Liu. 2014.
\newblock Topic modeling using topics from many domains, lifelong learning and
  big data.
\newblock In {\em ICML '14\/}. pages 703--711.

\bibitem[{Chen and Liu(2016)}]{ChenLiu2016}
Zhiyuan Chen and Bing Liu. 2016.
\newblock {\em Lifelong Machine Learning\/}.
\newblock Morgan {\&} Claypool Publishers.

\bibitem[{Chen et~al.(2015)Chen, Ma, and Liu}]{chen2015lifelong}
Zhiyuan Chen, Nianzu Ma, and Bing Liu. 2015.
\newblock Lifelong learning for sentiment classification.
\newblock {\em Volume 2: Short Papers\/} page 750.

\bibitem[{Chen et~al.(2014)Chen, Mukherjee, and Liu}]{ChenZhiyuan2014ACL}
Zhiyuan Chen, Arjun Mukherjee, and Bing Liu. 2014.
\newblock Aspect extraction with automated prior knowledge learning.
\newblock In {\em ACL '14\/}. pages 347--358.

\bibitem[{Choi and Cardie(2010)}]{Choi2010}
Yejin Choi and Claire Cardie. 2010.
\newblock Hierarchical sequential learning for extracting opinions and their
  attributes.
\newblock In {\em ACL '10\/}. pages 269--274.

\bibitem[{Fang and Huang(2012)}]{FangHuang2012ACL}
Lei Fang and Minlie Huang. 2012.
\newblock Fine granular aspect analysis using latent structural models.
\newblock In {\em ACL '12\/}. pages 333--337.

\bibitem[{Hu and Liu(2004)}]{HuL2004}
Minqing Hu and Bing Liu. 2004.
\newblock Mining and summarizing customer reviews.
\newblock In {\em KDD '04\/}. pages 168--177.

\bibitem[{Jakob and Gurevych(2010)}]{Jakob2010}
Niklas Jakob and Iryna Gurevych. 2010.
\newblock Extracting opinion targets in a single- and cross-domain setting with
  conditional random fields.
\newblock In {\em EMNLP '10\/}. pages 1035--1045.

\bibitem[{Jo and Oh(2011)}]{Jo2011}
Yohan Jo and Alice~H. Oh. 2011.
\newblock Aspect and sentiment unification model for online review analysis.
\newblock In {\em WSDM '11\/}. pages 815--824.

\bibitem[{Johansson and Moschitti(2010)}]{Johansson2010}
Richard Johansson and Alessandro Moschitti. 2010.
\newblock Syntactic and semantic structure for opinion expression detection.
\newblock In {\em Proceedings of the Fourteenth Conference on Computational
  Natural Language Learning\/}. pages 67--76.

\bibitem[{Lafferty et~al.(2001)Lafferty, McCallum, and
  Pereira}]{Lafferty2001conditional}
John Lafferty, Andrew McCallum, and Fernando~CN Pereira. 2001.
\newblock Conditional random fields: Probabilistic models for segmenting and
  labeling sequence data.
\newblock In {\em ICML '01\/}. pages 282--289.

\bibitem[{Li et~al.(2010)Li, Huang, and Zhu}]{LiHuangZhu2010}
Fangtao Li, Minlie Huang, and Xiaoyan Zhu. 2010.
\newblock Sentiment analysis with global topics and local dependency.
\newblock In {\em AAAI '10\/}. pages 1371--1376.

\bibitem[{Lin and He(2009)}]{Lin2009}
Chenghua Lin and Yulan He. 2009.
\newblock Joint sentiment/topic model for sentiment analysis.
\newblock In {\em CIKM '09\/}. pages 375--384.

\bibitem[{Liu(2012)}]{Liu2012}
Bing Liu. 2012.
\newblock {\em Sentiment Analysis and Opinion Mining\/}.
\newblock Morgan {\&} Claypool Publishers.

\bibitem[{Liu et~al.(2013)Liu, Xu, Liu, and Zhao}]{KangLiu2013IJCAI}
Kang Liu, Liheng Xu, Yang Liu, and Jun Zhao. 2013.
\newblock Opinion target extraction using partially-supervised word alignment
  model.
\newblock In {\em IJCAI '13\/}. pages 2134--2140.

\bibitem[{Liu et~al.(2016)Liu, Liu, Zhang, Kim, and Gao}]{liu2016improving}
Qian Liu, Bing Liu, Yuanlin Zhang, Doo~Soon Kim, and Zhiqiang Gao. 2016.
\newblock Improving opinion aspect extraction using semantic similarity and
  aspect associations.
\newblock In {\em Thirtieth AAAI Conference on Artificial Intelligence\/}.

\bibitem[{Mei et~al.(2007)Mei, Ling, Wondra, Su, and Zhai}]{MeiLWSZ2007}
Qiaozhu Mei, Xu~Ling, Matthew Wondra, Hang Su, and ChengXiang Zhai. 2007.
\newblock Topic sentiment mixture: Modeling facets and opinions in weblogs.
\newblock In {\em WWW '07\/}. pages 171--180.

\bibitem[{Mitchell et~al.(2013)Mitchell, Aguilar, Wilson, and
  Van~Durme}]{Mitchell-EtAl:2013:EMNLP}
Margaret Mitchell, Jacqui Aguilar, Theresa Wilson, and Benjamin Van~Durme.
  2013.
\newblock Open domain targeted sentiment.
\newblock In {\em ACL '13\/}. pages 1643--1654.

\bibitem[{Mitchell et~al.(2015)Mitchell, Cohen, Hruschka, Talukdar, Betteridge,
  Carlson, Dalvi, Gardner, Kisiel, Krishnamurthy, Lao, Mazaitis, Mohamed,
  Nakashole, Platanios, Ritter, Samadi, Settles, Wang, Wijaya, Gupta, Chen,
  Saparov, Greaves, and Welling}]{mitchell2015}
T~Mitchell, W~Cohen, E~Hruschka, P~Talukdar, J~Betteridge, A~Carlson, B~Dalvi,
  M~Gardner, B~Kisiel, J~Krishnamurthy, N~Lao, K~Mazaitis, T~Mohamed,
  N~Nakashole, E~Platanios, A~Ritter, M~Samadi, B~Settles, R~Wang, D~Wijaya,
  A~Gupta, X~Chen, A~Saparov, M~Greaves, and J~Welling. 2015.
\newblock Never-ending learning.
\newblock In {\em AAAI'2015\/}.

\bibitem[{Moghaddam and Ester(2011)}]{Moghaddam2011}
Samaneh Moghaddam and Martin Ester. 2011.
\newblock {ILDA}: interdependent lda model for learning latent aspects and
  their ratings from online product reviews.
\newblock In {\em SIGIR '11\/}. pages 665--674.

\bibitem[{Mukherjee and Liu(2012)}]{Mukherjee2012}
Arjun Mukherjee and Bing Liu. 2012.
\newblock Aspect extraction through semi-supervised modeling.
\newblock In {\em ACL '12\/}. volume~1, pages 339--348.

\bibitem[{Pan and Yang(2010)}]{pan2010survey}
Sinno~Jialin Pan and Qiang Yang. 2010.
\newblock A survey on transfer learning.
\newblock {\em IEEE Transactions on knowledge and data engineering\/}
  22(10):1345--1359.

\bibitem[{Popescu and Etzioni(2005)}]{PopescuNE2005}
Ana-Maria Popescu and Oren Etzioni. 2005.
\newblock Extracting product features and opinions from reviews.
\newblock In {\em HLT-EMNLP '05\/}. pages 339--346.

\bibitem[{Poria et~al.(2014)Poria, Cambria, Ku, Gui, and
  Gelbukh}]{poria2014rule}
Soujanya Poria, Erik Cambria, Lun-Wei Ku, Chen Gui, and Alexander Gelbukh.
  2014.
\newblock A rule-based approach to aspect extraction from product reviews.
\newblock In {\em SocialNLP '14\/}. pages 28--37.

\bibitem[{Qiu et~al.(2011)Qiu, Liu, Bu, and Chen}]{QiuLBC2011}
Guang Qiu, Bing Liu, Jiajun Bu, and Chun Chen. 2011.
\newblock Opinion word expansion and target extraction through double
  propagation.
\newblock {\em Computational Linguistics\/} 37(1):9--27.

\bibitem[{Ruvolo and Eaton(2013)}]{ruvolo2013ella}
Paul Ruvolo and Eric Eaton. 2013.
\newblock Ella: An efficient lifelong learning algorithm.
\newblock {\em ICML (1)\/} 28:507--515.

\bibitem[{Shu et~al.(2016)Shu, Liu, Xu, and Kim}]{shu2016lifelong}
Lei Shu, Bing Liu, Hu~Xu, and Annice Kim. 2016.
\newblock Lifelong-rl: Lifelong relaxation labeling for separating entities and
  aspects in opinion targets.
\newblock In {\em Proceedings of the Conference on Empirical Methods in Natural
  Language Processing (EMNLP)\/}.

\bibitem[{Silver et~al.(2013)Silver, Yang, and Li}]{silver2013lifelong}
Daniel~L Silver, Qiang Yang, and Lianghao Li. 2013.
\newblock Lifelong machine learning systems: Beyond learning algorithms.
\newblock In {\em AAAI Spring Symposium: Lifelong Machine Learning\/}.
  Citeseer, pages 49--55.

\bibitem[{Thrun(1998)}]{thrun1998lifelong}
Sebastian Thrun. 1998.
\newblock Lifelong learning algorithms.
\newblock In {\em Learning to learn\/}, Springer, pages 181--209.

\bibitem[{Titov and McDonald(2008)}]{TitovM2008}
Ivan Titov and Ryan McDonald. 2008.
\newblock A joint model of text and aspect ratings for sentiment summarization.
\newblock In {\em ACL '08: HLT\/}. pages 308--316.

\bibitem[{Wang and Wang(2008)}]{WangBo2008}
Bo~Wang and Houfeng Wang. 2008.
\newblock Bootstrapping both product features and opinion words from chinese
  customer reviews with cross-inducing.
\newblock In {\em IJCNLP '08\/}. pages 289--295.

\bibitem[{Wang et~al.(2010)Wang, Lu, and Zhai}]{Wang2010}
Hongning Wang, Yue Lu, and Chengxiang Zhai. 2010.
\newblock Latent aspect rating analysis on review text data: A rating
  regression approach.
\newblock In {\em KDD '10\/}. pages 783--792.

\bibitem[{Wang et~al.(2016)Wang, Chen, and Liu}]{WangWWW2016}
Shuai Wang, Zhiyuan Chen, and Bing Liu. 2016.
\newblock Mining aspect-speciﬁc opinion using a holistic lifelong topic
  model.
\newblock In {\em WWW '16\/}.

\bibitem[{Wu et~al.(2009)Wu, Zhang, Huang, and Wu}]{WuZHW2009}
Yuanbin Wu, Qi~Zhang, Xuanjing Huang, and Lide Wu. 2009.
\newblock Phrase dependency parsing for opinion mining.
\newblock In {\em EMNLP '09\/}. pages 1533--1541.

\bibitem[{Zhang et~al.(2010)Zhang, Liu, Lim, and O'Brien-Strain}]{Zhang2010}
Lei Zhang, Bing Liu, Suk~Hwan Lim, and Eamonn O'Brien-Strain. 2010.
\newblock Extracting and ranking product features in opinion documents.
\newblock In {\em COLING '10: Posters\/}. pages 1462--1470.

\bibitem[{Zhao et~al.(2010)Zhao, Jiang, Yan, and Li}]{ZhaoJiang2010}
Wayne~Xin Zhao, Jing Jiang, Hongfei Yan, and Xiaoming Li. 2010.
\newblock Jointly modeling aspects and opinions with a maxent-lda hybrid.
\newblock In {\em EMNLP '10\/}. pages 56--65.

\bibitem[{Zhao et~al.(2015)Zhao, Qin, and Liu}]{zhao2015creating}
Yanyan Zhao, Bing Qin, and Ting Liu. 2015.
\newblock Creating a fine-grained corpus for chinese sentiment analysis.
\newblock {\em IEEE Intelligent Systems\/} 30(1):36--43.

\bibitem[{Zhou et~al.(2013)Zhou, Wan, and Xiao}]{Zhou-wan-xiao:2013:EMNLP}
Xinjie Zhou, Xiaojun Wan, and Jianguo Xiao. 2013.
\newblock Collective opinion target extraction in {Chinese} microblogs.
\newblock In {\em EMNLP '13\/}. pages 1840--1850.

\bibitem[{Zhu et~al.(2009)Zhu, Wang, Tsou, and Zhu}]{Zhu2009}
Jingbo Zhu, Huizhen Wang, Benjamin~K. Tsou, and Muhua Zhu. 2009.
\newblock Multi-aspect opinion polling from textual reviews.
\newblock In {\em CIKM '09\/}. pages 1799--1802.

\bibitem[{Zhu(2005)}]{zhu2005semi}
Xiaojin Zhu. 2005.
\newblock Semi-supervised learning literature survey.
\newblock Technical Report 1530, Computer Sciences, University of
  Wisconsin-Madison.

\bibitem[{Zhuang et~al.(2006)Zhuang, Jing, and Zhu}]{ZhuangJZ2006}
Li~Zhuang, Feng Jing, and Xiao-Yan Zhu. 2006.
\newblock Movie review mining and summarization.
\newblock In {\em CIKM '06\/}. pages 43--50.

\end{thebibliography}

\documentclass[11pt,a4paper]{article}
\usepackage[hyperref]{acl2017}
\usepackage{times}
\usepackage{latexsym}
\usepackage{helvet}
\usepackage{courier}
\usepackage{amsmath}
\usepackage{graphicx}
\usepackage{fancyhdr}
\usepackage{adjustbox}
\usepackage{multirow}
\usepackage{float}
\usepackage{textcomp}
\usepackage{algorithm,algorithmic}
\usepackage{amssymb,amsthm}
\usepackage{bbm}
\usepackage{url}

\aclfinalcopy 
\def\aclpaperid{3125} 

\theoremstyle{definition}

\usepackage[belowskip=-15pt,aboveskip=0pt]{caption}

\makeatletter
\newcommand{\ALOOP}[1]{\ALC@it\algorithmicloop\ #1%
    \begin{ALC@loop}}
    \newcommand{\ENDALOOP}{\end{ALC@loop}\ALC@it\algorithmicendloop}

\newcommand{\algorithmicbreak}{\textbf{break}}
\newcommand{\BREAK}{\STATE \algorithmicbreak}
\makeatother

\title{Lifelong Learning CRF for Supervised Aspect Extraction}

\author{Lei Shu, Hu Xu, Bing Liu\\
    Department of Computer Science\\
    University of Illinois at Chicago\\
    \{lshu3, hxu48, liub\}@uic.edu
}

\date{}

\begin{document}

\maketitle
\begin{abstract}
This paper makes a focused contribution to supervised aspect extraction. It shows that if the system has performed aspect extraction from many past domains and retained their results as knowledge, Conditional Random Fields (CRF) can leverage this knowledge in a lifelong learning manner to extract in a new domain markedly better than the traditional CRF without using this prior knowledge. The key innovation is that even after CRF training, the model can still improve its extraction with experiences in its applications. 

\end{abstract}

\section{Introduction}
\label{sec:introduction}
Aspect extraction is a key task of opinion mining \cite{Liu2012}. It extracts opinion targets from opinion text. For example, from the sentence ``\emph{The battery is good}'', it aims to extract ``battery'', which is a product feature, also called an {\em aspect}. 


Aspect extraction 
is commonly done using a supervised or an unsupervised approach. The unsupervised approach includes methods such as frequent pattern mining~\cite{HuL2004,PopescuNE2005,Zhu2009}, syntactic rules-based extraction \cite{ZhuangJZ2006,WangBo2008,WuZHW2009,Zhang2010,QiuLBC2011,poria2014rule}, topic modeling \cite{MeiLWSZ2007,TitovM2008,LiHuangZhu2010,Brody2010,Wang2010,Moghaddam2011,Mukherjee2012,Lin2009,ZhaoJiang2010,Jo2011,FangHuang2012ACL,WangWWW2016}, word alignment \cite{KangLiu2013IJCAI}, label propagation \cite{Zhou-wan-xiao:2013:EMNLP,shu2016lifelong}, and others~\cite{zhao2015creating}.

This paper focuses on the supervised approach \cite{Jakob2010,Choi2010,Mitchell-EtAl:2013:EMNLP} using Conditional Random Fields (CRF) \cite{Lafferty2001conditional}. It shows that the results of CRF can be significantly improved by leveraging some prior knowledge automatically mined from the extraction results of previous domains, including domains without labeled data. 
The improvement is possible because although every product (domain) is different, there is a fair amount of aspects sharing across domains~\cite{ChenLiu2014ICML}. For example, every review domain has the aspect \textit{price} and reviews of many products have the aspect \textit{battery life} or \textit{screen}. Those shared aspects may not appear in the training data but appear in unlabeled data and the test data.
We can exploit such sharing to help CRF perform much better.

Due to leveraging the knowledge gained from the past to help the new domain extraction, we are using the idea of {\em lifelong machine learning} (LML)~\cite{ChenLiu2016,thrun1998lifelong,silver2013lifelong}, which is a continuous learning paradigm that retains the knowledge learned in the past and uses it to help future learning and problem solving with possible adaptations. 

The setting of the proposed approach L-CRF ({\em Lifelong CRF}) is as follows: A CRF model $M$ has been trained with a labeled training review dataset. At a particular point in time, $M$ has extracted aspects from data in $n$ previous domains $D_1, \dots, D_n$ (which are unlabeled) and the extracted sets of aspects are $A_1, \dots, A_n$. Now, the system is faced with a new domain data $D_{n+1}$.  $M$ can leverage some {\em reliable prior knowledge} in $A_1, \dots, A_n$ to make a better extraction from $D_{n+1}$ than without leveraging this prior knowledge. 

The key innovation of L-CRF is that even after supervised training, the model can still improve its extraction in testing or its applications with experiences. Note that L-CRF is different from semi-supervised learning \cite{zhu2005semi} as the $n$ previous (unlabeled) domain data used in extraction are not used or not available during model training.

There are prior LML works for aspect extraction \cite{ChenZhiyuan2014ACL,liu2016improving}, but they were all unsupervised methods. Supervised LML methods exist \cite{chen2015lifelong,ruvolo2013ella}, but they are for classification rather than for sequence learning or labeling like CRF. A semi-supervised LML method is used in NELL\cite{mitchell2015}, but it is heuristic pattern-based. It doesn't use sequence learning and is not for aspect extraction. LML is related to transfer learning and multi-task learning~\cite{pan2010survey}, but they are also quite different (see~\cite{ChenLiu2016} for details).


To the best of our knowledge, this is the first paper that uses LML to help a supervised extraction method to markedly improve its results.  

\section{Conditional Random Fields}
CRF learns from an observation sequence $\mathbf{x}$ to estimate a label sequence $\mathbf{y}$: $p(\mathbf{y}|\mathbf{x}; \boldsymbol{\theta})$, where $\boldsymbol{\theta}$ is a set of weights. 
Let $l$ be the $l$-th position in the sequence. The core parts of CRF are a set of feature functions $\mathcal {F}=\{f_h(y_{l},y_{l-1},\mathbf{x}_l)\}_{h=1}^{H}$ and their corresponding weights $\boldsymbol{\theta}=\{\theta_h\}_{h=1}^H$. 

\noindent
{\bf Feature Functions}: We use two types of feature functions (FF). 
One is \emph{Label-Label (LL)} FF:
    \begin{equation}
    \label{ll feature function}
    \resizebox{0.85\hsize}{!}{
        $f_{ij}^{\textit{LL}}(y_l,y_{l-1})=\mathbbm{1}\{y_l=i\}\mathbbm{1}\{y_{l-1}=j\} ,\forall i,j \in \mathcal{Y},$
    }
    \end{equation}
where $\mathcal{Y}$ is the set of labels, and $\mathbbm{1}\{\cdot\} $ an indicator function.
The other is \emph{Label-Word (LW)} FF:
    \begin{equation}
    \label{lw feature function}
    \resizebox{0.85\hsize}{!}{
        $f_{iv}^{\textit{LW}}(y_l,\mathbf{x}_l) = \mathbbm{1}\{y_l=i\}\mathbbm{1}\{\mathbf{x}_l=v\} ,\forall i \in \mathcal{Y}, 
        \forall v \in \mathcal{V},$
    }
    \end{equation}
where $\mathcal{V}$ is the vocabulary. This FF returns $1$ when the $l$-th word is $v$ and the $l$-th label is $v$'s specific label $i$; otherwise $0$. $\mathbf{x}_l$ is the current word, and is represented as a multi-dimensional vector. Each dimension in the vector is a feature of $\mathbf{x}_l$.  

Following the previous work in \cite{Jakob2010}, 
we use the feature set \{W, -1W, +1W, P, -1P, +1P, G\}, where W is the word and P is its POS-tag, -1W is the previous word, -1P is its POS-tag, +1W is the next word, +1P is its POS-tag, and G is the generalized dependency feature. 

Under the Label-Word FF type, we have two sub-types of FF: {\em Label-dimension} FF and {\em Label-G} FF. Label-dimension FF is for the first 6 features, and Label-G is for the G feature. 

The \emph{Label-dimension (L$d$)} FF is defined as
\begin{equation}
\label{ld feature function}
\resizebox{0.85\hsize}{!}{
    $f_{iv^d}^{\textit{L}d}(y_l,\mathbf{x}_l) = \mathbbm{1}\{y_l=i\}\mathbbm{1}\{\mathbf{x}_l^d=v^d\}, \forall i \in \mathcal{Y}, 
    \forall v^d \in \mathcal{V}^d ,$
}
\end{equation}
where $\mathcal{V}^d$ is the set of observed values in feature $d \in \{W, -1W, +1W, P, -1P, +1P\}$ and we call $\mathcal{V}^d$ feature $d$'s feature values. Eq. (\ref{ld feature function}) is a FF that returns $1$ when $\mathbf{x}_l$'s feature $d$ equals to the feature value $v^d$ and the variable $y_l$ ($l$th label) equals to the label value $i$; otherwise 0.

We describe G  and its feature function next, which also holds the key to the proposed L-CRF. 
        
\begin{table*}[ht]
    \centering
    \scalebox{0.75}{
    \begin{tabular}{ c | c| l }
        \hline
        Index & Word & Dependency Relations  \\
        \hline
        1 & The & \{\textit{(det, battery, 2, NN , The, 1, DT)} \}\\
        \hline
        2 & battery &\{\textit{(nsubj, great, 7, JJ , battery, 2, NN), (det, battery, 2, NN , The, 1, DT), (nmod, battery, 2, NN, camera, 5, NN)} \} \\
        \hline
        3 & of &\{\textit{(case, camera, 5, NN, of, 3, IN)} \}\\
        \hline
        4 & this&\{\textit{(det, camera, 5, NN, this, 4, DT)} \}\\
        \hline
        5 & camera&\{\textit{(case, camera, 5, NN, of, 3, IN), (det, camera, 5, NN, this, 4, DT), (nmod, battery, 2, NN, camera, 5, NN)} \} \\
        \hline
        6 & is&\{\textit{(cop, great, 7, JJ , is, 6, VBZ)} \}\\
        \hline
        7 & great&\{\textit{(root, ROOT, 0, VBZ, great, 7, JJ), (nsubj, great, 7, JJ , battery, 2, NN), (cop, great, 7, JJ , is, 6, VBZ)} \} \\
        \hline 
    \end{tabular}
    }
    \caption{Dependency relations parsed from ``The battery of this camera is great''}
    \label{table:dr1}
\end{table*}
\section{General Dependency Feature (G)}
Feature G uses generalized dependency relations. What is interesting about this feature is that it enables L-CRF to use past knowledge in its sequence prediction at the test time in order to perform much better. This will become clear shortly. 
This feature 
takes a \emph{dependency pattern} as its value, which is generalized from dependency relations.

The general dependency feature (G) of the variable $\mathbf{x}_l$ takes a set of feature values $\mathcal{V}^{\textit{G}}$. Each feature value $v^{\textit{G}}$ is a dependency pattern. The \emph{Label-G (LG)} FF is defined as:
    \begin{equation}
    \label{lg feature function}
    \resizebox{0.85\hsize}{!}{
        $f_{iv^{\text{G}}}^{\textit{LG}}(y_l,\mathbf{x}_l) = \mathbbm{1}\{y_l=i\}\mathbbm{1}\{\mathbf{x}_l^{\textit{G}}=v^{\textit{G}}\} ,\forall i \in \mathcal{Y}, 
        \forall v^{\text{G}}\in \mathcal{V}^{\text{G}}.$
    }
    \end{equation}
Such a FF returns $1$ when the dependency feature of the variable $\mathbf{x}_l$ equals to a dependency pattern $v^{\textit{G}}$ and the variable $y_l$ equals to the label value $i$.




\subsection{Dependency Relation}
Dependency relations have been shown useful in many sentiment analysis applications \cite{Johansson2010,Jakob2010}. A dependency relation \footnote{We obtain dependency relations using Stanford CoreNLP: http://stanfordnlp.github.io/CoreNLP/.} is a quintuple-tuple:  
$\textit{(type, gov, govpos, dep, deppos)},$
where $\textit{type}$ is the type of the dependency relation, \textit{gov} is the \emph{governor word}, \textit{govpos} is the POS tag of the governor word, \textit{dep} is the \emph{dependent word}, and \textit{deppos} is the POS tag of the dependent word. The $l$-th word can either be the governor word or the dependent word in a dependency relation.



\subsection{Dependency Pattern}
We generalize dependency relations into {\em dependency patterns} using the following steps: 

\begin{enumerate}
    \item For each dependency relation, replace the current word (governor word or dependent word) and its POS tag with a wildcard since we already have the word (W) and the POS tag (P) features.

    \item Replace the context word (the word other than the $l$-th word) in each dependency relation with a knowledge label to form a more general dependency pattern. Let the set of aspects annotated in the training data be $K^t$. If the context word in the dependency relation appears in $K^t$, we replace it with a knowledge label `A' (aspect); otherwise `O' (other).

\end{enumerate}

For example, we work on the sentence ``The battery of this camera is great.'' The dependency relations are given in Table \ref{table:dr1}. Assume the current word is ``battery,'' and ``camera'' is annotated as an aspect. The original dependency relation between ``camera'' and ``battery'' produced by a parser is (nmod, battery, NN, camera, NN). Note that we do not use the word positions in the relations in Table \ref{table:dr1}. Since the current word's information (the word itself and its POS-tag) in the dependency relation is redundant, we replace it with a wild-card. The relation becomes (nmod, *, camera, NN). Secondly, since ``camera'' is in $K^t$, we replace ``camera'' with a general label `A'. The final dependency pattern becomes (nmod,*, A, NN).


We now explain why dependency patterns can enable a CRF model to leverage the past knowledge. The key is the knowledge label `A' above, which indicates a likely aspect. Recall that our problem setting is that when we need to extract from the new domain $D_{n+1}$ using a trained CRF model $M$, we have already extracted from many previous domains $D_1, \dots, D_n$ and retained their extracted sets of aspects $A_1, \dots, A_n$. Then, we can mine reliable aspects from $A_1, \dots, A_n$ and add them in $K^t$, which enables many knowledge labels in the dependency patterns of the new data $A_{n+1}$ due to sharing of aspects across domains. This enriches the dependency pattern features, which consequently allows more aspects to be extracted from the new domain $D_{n+1}$.

\begin{algorithm}
    \caption{Lifelong Extraction of L-CRF}\label{algo:pred}
    \begin{algorithmic}[1]
        \STATE $K_p \leftarrow \emptyset$
        \ALOOP{} 
        \STATE $F \leftarrow \text{FeatureGeneration}(D_{n+1}, K)$ \label{test:1}
        \STATE $A_{n+1} \leftarrow \text{Apply-CRF-Model}(M, F)$  \label{test:2}
        \STATE $S \leftarrow S \cup \{A_{n+1}\}$
        \STATE $K_{n+1} \leftarrow \text{Frequent-Aspects-Mining}(S,\lambda)$
        \IF {$K_p = K_{n+1}$}
        \BREAK
        \ELSE
        \STATE $K \leftarrow K^t \cup K_{n+1}$\label{test:3}
        \STATE $K_p \leftarrow K_{n+1}$\label{test:4}
        \STATE $S \leftarrow S - \{A_{n+1}\}$
        \ENDIF
        \ENDALOOP
    \end{algorithmic}
\end{algorithm}
\section{The Proposed L-CRF Algorithm}

We now present the L-CRF algorithm. As the dependency patterns for the general dependency feature do not use any actual words and they can also use the prior knowledge, they are quite powerful for cross-domain extraction (the test domain is not used in training). 

Let $K$ be a set of {\em reliable aspects} mined from the aspects extracted in past domain datasets using the CRF model $M$. Note that we assume that $M$ has already been trained using some labeled training data $D^t$. Initially, $K$ is $K^t$ (the set of all annotated aspects in the training data $D^t$). The more domains $M$ has worked on, the more aspects it extracts, and the larger the set $K$ gets. When faced with a new domain $D_{n+1}$, $K$ allows the general dependency feature to generate more dependency patterns related to aspects due to more knowledge labels `A' as we explained in the previous section. Consequently, CRF has more informed features to produce better extraction results. 

L-CRF works in two phases: {\em training phase} and {\em lifelong extraction phase}. The training phase trains a CRF model $M$ using the training data $D^t$, which is the same as normal CRF training, and will not be discussed further. In the lifelong extraction phase, $M$ is used to extract aspects from coming domains ($M$ does not change and the domain data are unlabeled). All the results from the domains are retained in past aspect store $S$. At a particular time, it is assumed $M$ has been applied to $n$ past domains, and is now faced with the $n+1$ domain. L-CRF uses $M$ and reliable aspects (denoted $K_{n+1}$) mined from $S$ and $K^t$ ($K=K^t \cup K_{n+1}$) to extract from $D_{n+1}$.  Note that aspects $K_t$ from the training data are considered always reliable as they are manually labeled, thus a subset of $K$. We cannot use all extracted aspects from past domains as reliable aspects due to many extraction errors. But those aspects that appear in multiple past domains are more likely to be correct. Thus $K$ contains those frequent aspects in $S$. The lifelong extraction phase is in Algorithm \ref{algo:pred}.

\textbf{Lifelong Extraction Phase}: 
Algorithm \ref{algo:pred} performs extraction on $D_{n+1}$ iteratively. 
\begin{enumerate}
    \item It generates features ($F$) on the data $D_{n+1}$ (line \ref{test:1}), and applies the CRF model $M$ on $F$ to produce a set of aspects $A_{n+1}$ (line \ref{test:2}). 
    \item $A_{n+1}$ is added to $S$, the past aspect store. From $S$, we mine a set of frequent aspects $K_{n+1}$. The frequency threshold is $\lambda$.
    \item If $K_{n+1}$ is the same as $K_p$ from the previous iteration, the algorithm exits as no new aspects can be found. We use an iterative process because each extraction gives new results, which may increase the size of $K$, the reliable past aspects or past knowledge. The increased $K$ may produce more dependency patterns, which can enable more extractions. 
    \item Else: some additional reliable aspects are found. $M$ may extract additional aspects in the next iteration. Lines \ref{test:3} and \ref{test:4} update the two sets for the next iteration.
\end{enumerate}


     \begin{table}[t]
         \centering
            \scalebox{0.85}{
             \begin{tabular}{c|c|c|c}
                 \hline
                 {\bf Domain}&{\bf \# Sent.}& {\bf \# Asp.}& {\bf \# non-asp. words}  \\\hline
                 Computer     &536    &1173    &7675\\\hline
                 Camera     &609    &1640    &9849\\\hline
                 Router     &509    &1239    &7264\\\hline
                 Phone         &497    &980    &7478\\\hline
                 Speaker      &510    &1299    &7546\\\hline
                 DVD Player &506    &928    &7552\\\hline
                 Mp3 Player &505    &1180     &7607\\\hline
             \end{tabular}
             }
        \caption{{Annotation details of the datasets}}\label{tab:annotation} 
     \end{table}

\begin{table*}[ht]
    \centering
    \scalebox{0.95}{
    \begin{tabular}{c|c|c c c|c c c|c c c }
            \hline    
            \multicolumn{11}{c}{\textbf{Cross-Domain}} \\\hline  
            \multirow{2}{*}{Training}&\multirow{2}{*}{Testing}&\multicolumn{3}{ |c| }{CRF}& \multicolumn{3}{ |c| }{CRF+R}& \multicolumn{3}{ |c }{L-CRF} \\

            &&$\mathcal{P}$& $\mathcal{R}$&$\mathcal{F}_1$&$\mathcal{P}$& $\mathcal{R}$&$\mathcal{F}_1$
            &$\mathcal{P}$& $\mathcal{R}$&$\mathcal{F}_1$\\\hline

            $-$Computer&Computer&86.6&51.4&64.5&23.2&90.4&37.0&82.2&62.7&71.1\\\hline
            $-$Camera&Camera&84.3&48.3&61.4&21.8&86.8&34.9&81.9&60.6&69.6\\\hline
            $-$Router&Router&86.3&48.3&61.9&24.8&92.6&39.2&82.8&60.8&70.1\\\hline
            $-$Phone&Phone&72.5&50.6&59.6&20.8&81.2&33.1&70.1&59.5&64.4\\\hline
            $-$Speaker&Speaker&87.3&60.6&71.6&22.4&91.2&35.9&84.5&71.5&77.4\\\hline
            $-$DVDplayer&DVDplayer&72.7&63.2&67.6&16.4&90.7&27.7&69.7&71.5&70.6\\\hline
            $-$Mp3player&Mp3player&87.5&49.4&63.2&20.6&91.9&33.7&84.1&60.7&70.5\\\hline\hline
            &\textbf{Average}&82.5&53.1&64.3&21.4&89.3&34.5&79.3&63.9&70.5\\\hline 
  
            \multicolumn{11}{c}{\textbf{In-Domain}}\\\hline
            $-$Computer&$-$Computer&84.0&71.4&77.2&23.2&93.9&37.3&81.6&75.8&78.6\\\hline
            $-$Camera&$-$Camera&83.7&70.3&76.4&20.8&93.7&34.1&80.7&75.4&77.9\\\hline
            $-$Router&$-$Router&85.3&71.8&78.0&22.8&93.9&36.8&82.6&76.2&79.3\\\hline
            $-$Phone&$-$Phone&85.0&71.1&77.5&25.1&93.7&39.6&82.9&74.7&78.6\\\hline
            $-$Speaker&$-$Speaker&83.8&70.3&76.5&20.1&94.3&33.2&80.1&75.8&77.9\\\hline
            $-$DVDplayer&$-$DVDplayer&85.0&72.2&78.1&20.9&94.2&34.3&81.6&76.7&79.1\\\hline
            $-$Mp3player&$-$Mp3player&83.2&72.6&77.5&20.4&94.5&33.5&79.8&77.7&78.7\\\hline\hline
            &\textbf{Average}&84.3&71.4&77.3&21.9&94.0&35.5&81.3&76.0&78.6\\\hline             
        \end{tabular}
        }
    \caption{Aspect extraction results in precision, recall and F$_1$ score: Cross-Domain and In-Domain ($-$X means all except domain X)} 
    \label{tab:result} 
\end{table*}

\section{Experiments} 
    We now evaluate the proposed L-CRF method and compare with baselines.  
    
\subsection{Evaluation Datasets}
    We use two types of data for our experiments. The first type consists of seven (7) annotated benchmark review datasets from 7 domains (types of products). Since they are annotated, they are used in training and testing. The first 4 datasets are from~\cite{HuL2004}, which actually has 5 datasets from 4 domains. Since we are mainly interested in results at the domain level, we did not use one of the domain-repeated datasets. The last 3 datasets of three domains (products) are from~\cite{liu2016improving}. These datasets are used to make up our CRF training data $D^t$ and test data $D_{n+1}$. The annotation details are given in Table \ref{tab:annotation}. 
    

    
    The second type has 50 unlabeled review datasets from 50 domains or types of products~\cite{ChenLiu2014ICML}. Each dataset has 1000 reviews. They are used as the past domain data, i.e., $D_1, \dots, D_n$ ($n=50$). Since they are not labeled, they cannot be used for training or testing. 
    
    
    
\subsection{Baseline Methods} 
    We compare L-CRF with CRF. We will not compare with unsupervised methods, which have been shown improvable by lifelong learning~\cite{ChenZhiyuan2014ACL,liu2016improving}. The frequency threshold $\lambda$ in Algorithm 1 used in our experiment to judge which extracted aspects are considered reliable is empirically set to $2$.    
    
    {\bf CRF}: We use the linear chain CRF from  \footnote{https://github.com/huangzhengsjtu/pcrf/}. Note that CRF uses all features including dependency features as the proposed L-CRF but does not employ the 50 domains unlabeled data used for lifelong learning
    
    {\bf CRF+R}: It treats the reliable aspect set $K$ as a dictionary. It adds those reliable aspects in $K$ that are not extracted by CRF but are in the test data to the final results. We want to see whether incorporating $K$ into the CRF extraction through dependency patterns in L-CRF is actually needed.     

We do not compare with domain adaptation or transfer learning because domain adaption basically uses the source domain labeled data to help learning in the target domain with few or no labeled data. Our 50 domains used in lifelong learning have no labels. So they cannot help in transfer learning. Although in transfer learning, the target domain usually has a large quantity of unlabeled data, but the 50 domains are not used as the target domains in our experiments. 
    
    
\subsection{Experiment Setting} 

To compare the systems using the same training and test data, for each dataset we use 200 sentences for training and 200 sentences for testing to avoid bias towards any dataset or domain because we will combine multiple domain datasets for CRF training. We conducted both cross-domain and in-domain tests. Our problem setting is cross-domain. In-domain is used for completeness. In both cases, we assume that extraction has been done for the 50 domains. 

    {\bf Cross-domain experiments}: We combine 6 labeled domain datasets for training (1200 sentences) and test on the 7th domain (not used in training). This gives us 7 {\em cross-domain} results. This set of tests is particularly interesting as it is desirable to have the trained model used in cross-domain situations to save manual labeling effort. 
    
    {\bf In-domain experiments}: We train and test on the same 6 domains (1200 sentences for training and 1200 sentences for testing). This also gives us 7 {\em in-domain} results. 

    {\bf Evaluating Measures}: 
    We use the popular precision $\mathcal{P}$, recall $\mathcal{R}$, and $\mathcal{F}_1$-score. 
    
    \subsection{Results and Analysis}
    
    All the experiment results are given in Table \ref{tab:result}. 
    
       {\bf Cross-domain}: Each $-$X in column 1 means that domain X is not used in training. X in column 2 means that domain X is used in testing. We can see that L-CRF is markedly better than CRF and CRF+R in $\mathcal{F}_1$. CRF+R is very poor due to poor precisions, which shows treating the reliable aspects set $K$ as a dictionary isn't a good idea. 
        
        
       {\bf In-domain}: $-$X in training and test columns means that the other 6 domains are used in both training and testing (thus in-domain). We again see that L-CRF is consistently better than CRF and CRF+R in $\mathcal{F}_1$. The amount of gain is smaller. This is expected because most aspects appeared in training probably also appear in the test data as they are reviews from the same 6 products.
        
        
    
    \section{Conclusion}
    
    This paper proposed a lifelong learning method to enable CRF to leverage the knowledge gained from extraction results of previous domains (unlabeled) to improve its extraction. Experimental results showed the effectiveness of L-CRF. The current approach does not change the CRF model itself. In our future work, we plan to modify CRF so that it can consider previous extraction results as well as the knowledge in previous CRF models. 
    
\section*{Acknowledgments} 
This work was supported in part by grants from National Science Foundation (NSF) under grant no. IIS-1407927 and IIS-1650900.

    \bibliography{acl2017}
    \bibliographystyle{acl_natbib}

\end{document}